OpenStudy (idealist10):

A radioactive material loses 25% of its mass in 10 minutes. Find its half-life.

OpenStudy (idealist10):

@hartnn @satellite73 @Compassionate

OpenStudy (anonymous):

\[.75^{\frac{t}{10}}=.5\] solve for \(t\)

OpenStudy (anonymous):

\[\frac{t}{10}=\frac{\ln(.5)}{\ln(.75)}\] by the change of base formula

OpenStudy (idealist10):

Thank you so much!